if are the roots of the equation . Then what is the value of .

 if  are the roots of the equation . Then what is the value of .

Grade:10

4 Answers

Sathya
12 years ago

      

 

    

 

         

Mir Noaman Ali IIT-Roorkee
12 years ago

Dear Sathyaram

 a (alpha) b (beta) a+b =1 so (a+b)^2012=1 now (a-b)^2=(a+b)^2-2ab=1-2=-1 so a-b whole square =-1 hence a-b=i (iota),(a-b)^2012=i^2012=1 so answer is 1+1=2 , Please Approve my answer if you have understood the solution 

Regards

Noaman Ali

Swapnil Saxena
12 years ago

the answer for the question is 1+(3)1006

Vaishal Shah
12 years ago

1+3^1006